OpenStudy (anonymous):

Matthew is waterskiing. As the boat starts moving, he is at an angle of 8.0° to the right of the boat. The boat applies 250 newtons over the first 50 meters, while the angle stays at roughly 8.0°. What work has the boat done on Matthew

OpenStudy (anonymous):

 workdone=Fd*cos(theta) =250*50*cos(8) =12500*cos(8)=12378.35joule